in an effort to control vegetation overgrowth, 75 rabbits are released in an isolated area free of predators. After 3 year, it is estimated that the rabbit population has increased to 760. Assuming exponential population growth, what will the population be after another 6 months? Round to the nearest rabbit.
Answer by stanbon(75887)

in an effort to control vegetation overgrowth, 75 rabbits are released in an isolated area free of predators. After 3 year, it is estimated that the rabbit population has increased to 760. Assuming exponential population growth, what will the population be after another 6 months? Round to the nearest rabbit.
------
Growth model:: y = ab^t
----
Using the data you get:
760 = 75b^3
b = 2.164
--------------
Equation:
g(t) = 75*2.164^t
-----
g(6) = 75*2.164^6 = 7701 when rounded down
---------------------------
Cheers,
Stan H.
------------------